Video Concepts, Inc. (VCI) markets independently produced films on DVDs through a variety of video retail outlets. VCI is deciding whether they should obtain the rights to a film titled Burnt Orange Forever. VCI estimated the total market for the film will be 100,000 units.

Other data available:

Distribution rights for the film: $100,000.00

DVD Label design: $5,000.00

DVD Package design: $10,000.00

Advertising: $35,000.00

Reproduction of copies: $4,000.00 per 1,000 units

Manufacture of packaging: $750.00 per 1,000 units

Royalties: $500.00 per 1,000 units

DVD Shipping Costs: $1 per unit

VCI's suggested retail price for the film is $20/unit and the retailer's margin is 50%. (There are no other channel intermediaries such as wholesalers to consider.)

1. What is VCI's $ contribution margin per DVD unit sold?

2. What is VCI's % contribution margin per unit sold?

3. How many units will VCI have to sell to break even?

4. If VCI sells 100K units, what will their $ profit be?

5. VCI is considering a shift to digital distribution, replacing DVD production with digital downloads. Calculate the breakeven units (downloads) required given the following data:

  • Lower retail price of $15.00
  • Digital distribution retailers (iTunes, Amazon) will require a 25% margin
  • VCI will sill need to pay distribution rights for the film, as well as royalites for each download.
  • VCI will still need to design a digital "poster" for the film, costing $5,000.
  • VCI will not need to produce DVDs, design/manufacture packaging, or ship DVDs
  • As there will not be brick-and-mortar retail outlets, a higher level of advertising will be necessary: $50,000.00
Answer #1

Total Fixed Cost = 100,000 + 5,000 + 10,000 + 35,000 = $150,000

Variable Cost per unit = (4000+ 750 + 500)/1000 + 1 =$ 6.25

Retail Price per unit = $ 20

Retailer margin = 50%

Total Variable Cost per unit = 6.25 + 50/100 *$20 = 6.25+ 10 = $ 16.25

1. $ contribution margin = Net Price per unit - Variable Cost per unit =$20 - $10 - $6.25 = $3.75/DVD unit sold

2. % contribution margin = 3.75/20 *100 = 18.75%

3. for break even(Let total sales =x)

=> Total Costs = Total Revenue

=> FC+VC = Sales Revenue => 150,000 + 16.25*x = 20*x

=> 3.75*x = 150,000

=> x= 40,000

4. IF VCI selles 100k units=> Their net sales revenue = 20* 100,000 = $ 2,000,000

Their total cost = 150,000 + 16.25* 100,000 = 1,775,000

Total Profit = Sales Revenue - Total Cost = 2,000,000 - 1,775,000 = $ 225,000